OpenStudy (anonymous):

factor plse 81r^6 - 3f^6 =

OpenStudy (anonymous):

3*(27r^6 - f^6)=3*(3^3 (r^2)^3 -(f^2)^3)= 3(3r^2 - f^2)(9r^4-3r^2*f^2 + f^4)

OpenStudy (anonymous):

inik can you be a little more detailed

OpenStudy (anonymous):

i'll try...

OpenStudy (anonymous):

remember, that: 27=3^3 r^6=(r^2)^3 f^6=(f^2)^3 and use formula for( a^3-b^3)

OpenStudy (anonymous):

\[=3*((3r ^{2})^{3}-(f ^{2})^{3})=3(3r ^{2}+f ^{2})*(3^{2}r ^{4}-3r ^{2}f ^{2}+f ^{4})\] is it better?

OpenStudy (anonymous):

shooot! it should be "-" : (3r^2 -f^2) and plus in second parenthesis... (see formula for a^3-b^3) please correct in your final answer

OpenStudy (anonymous):

should be: \[3(3r ^{2}-f ^{2})(9r ^{4}+3r ^{2}f ^{2}+f ^{4})\]